# Jina-ColBERT-v2: A General-Purpose Multilingual Late Interaction Retriever

### v3

@@ Abstract
Multi-vector dense models, such as ColBERT, have proven highly effective in information retrieval. ColBERT’s late interaction scoring approximates the joint query-document attention seen in cross-encoders while maintaining inference efficiency closer to traditional dense retrieval models, thanks to its bi-encoder architecture and recent optimizations in indexing and search. In this paper, we introduce a novel architecture and a training framework to support long context window and multilingual retrieval. Our new model, Jina-ColBERT-v2, demonstrates strong performance across a range of English and multilingual retrieval tasks,
@@ 1 Introduction
Neural retrieval has gained popularity in recent years following the arrival of capable pre-trained language models (PLMs) (Devlin et al., 2019; Liu et al., 2019; Clark et al., 2020). Two types of approaches have been employed to apply PLMs in retrieval tasks. Sparse neural retrieval systems, such as SPLADE (Formal et al., 2021), represent texts as weighted bags of words. In contrast, dense retrievers encode queries and documents as high-dimensional vectors, capturing relevance signals through spatial relationships extending beyond exact term matching.
Most dense retrievers encode a query or document as a single vector, commonly the result of mean-pooling or the [CLS]-embedding over the transformer’s final layer token embeddings. The relevance is calculated between the query vector and the document vector. In comparison, multi-vector retrievers like ColBERT (Khattab and Zaharia, 2020) generalize this embedding process to maintain a embedding for each token, computing relevance scores as a late interaction over token embeddings.
Storing embeddings for each token, ColBERT models preserve the information at the most fine-grained granularity. With the late interaction functioned at the query time, ColBERT achieves better in-domain performance and tends to be more robust out-of-domain than single-vector similarity (Khattab and Zaharia, 2020). However, the context length of ColBERT is limited to 512 tokens which restricted the information stored in a single chunk.
Another drawback is that ColBERT takes more time to calculate because it needs to combine token interactions into one final score based on the embeddings of each token. Although it works well with the existing vector similarity infrastructure that makes single-vector methods efficient, ColBERT requires more storage space for the embeddings. To make the ColBERT usable in practice, the output dimensionality is restricted to be much smaller than the single-vector models.
Moreover, ColBERTv2 (Santhanam et al., 2022) is trained only on English MSMARCO triplets (Bajaj et al., 2016) and has a monolingual BERT backbone and is incapable of multilingual retrieval. Some previous works have made to extend the model to multilingual retrieval tasks. Louis et al. (2024) does this by using parameter extensions for each additional language, and Lawrie et al. (2023) trains solely on machine-translated English MSMARCO data to get effective multilingual performance. Due to the lack of diversity in the training corpus, these approaches come with significant trade-offs in terms of model usability and training data diversity.
@@
  1. 1.
    Training with diverse weakly-supervised data: Equipped with rotary position embedding and trained with large-scale unlabeled text pairs from various corpus with weakly-supervised contrastive pretraining, our model is able to offer high-quality multi-vector embeddings for 8k-token-long text. A second-stage finetuning with labeled data using supervised distillation is used to further boost its performances.
  2. 2.
    General multilingual performance: We train with data from a variety of high- and low-resource languages using both labeled and unlabeled data, including machine-translated training data, and show that this improves even out-of-domain multilingual performance. The diversity of the training data provide good performance in general on the multilingual retrieval tasks.
  3. 3.
    Inference-agnostic efficiency: We introduce multiple sizes of linear projection heads, jointly trained using the non-weight tying variant of Matryoshka Representation Loss Kusupati et al. (2022), enabling the selection of token embedding size at inference time with minimal performance degradation. We demonstrate that reducing the embedding dimensionality from 128 to 64 only introduces insignificant performance drop. Additionally, our flash-attention optimized backbone, Jina-XLM-RoBERTa  provides further performance boosts during inference.
Our experimental results show competitive retrieval performance across both English and multilingual benchmarks. We also present controlled experiments demonstrating the benefit, or lack thereof, of the training modifications we consider in developing our training recipe.
@@ 2.3 Multilingual Retrieval
Owing to the quality of English-based pre-trained models (BERT) and annotated data (MSMARCO), many advances in neural retrieval have been in a monolingual English setting Karpukhin et al. (2020); Xiong et al. (2020); Khattab and Zaharia (2020). Researchers, however, have also made advances in non-English capabilities. On the modeling front, multilingual PLMs like mBERT (Devlin et al., 2019) and later XLM-RoBERTa (Conneau et al., 2020) have expanded pre-training to include text in up to 100 languages, including in cross-language contexts. For multilingual retrieval, there are two approaches: natural and translated. Datasets like Mr-Tydi and MIRACL (Zhang et al., 2021, 2023b) are built from human-generated and annotated queries, whereas mMARCO (Bonifacio et al., 2022) is a collection of machine-translated copies of MSMARCO which inherit their judgments from the original dataset. The former method tends to be of higher quality and lacks the subtle distributional/idiomatic errors, dubbed "translationese", that the latter sometimes exhibits. Naturally, however, human generation costs more per example.
@@ 4.1 Backbone Improvements
In order to support long context window, We replace the absolute positional embeddings with rotary positional embeddings (RoPE, Su et al. (2023)). To warm up its new positional embeddings, we continued pre-training the modified backbone with the same masked language modeling objective for 160,000 steps on the RefinedWeb dataset Penedo et al. (2023), a modern, high-quality corpus. During this pre-training phase, we set the maximum sequence length to 8,192 tokens with a rotary base of 10,000 and employed whole-word-masking Devlin et al. (2019), masking out 30% of the tokens. We call this modified language model Jina-XLM-RoBERTa.
@@ 4.2 Multiple Linear Heads
To reduce index sizes, ColBERT includes a linear head that projects its token embeddings from the hidden dimension of its language model down to a lower dimension ($768\to 128$). Notably, BGE-M3’s multi-vector retrieval does not take this step, keeping its token embeddings at a full 1024 dimensions.
To further reduce index size, we jointly train six linear heads with dimensions $d\in\{64,96,128,256,512,768\}$ using Matryoshka Representation Loss (MRL, Kusupati et al. (2022)). This allows users to choose greater or lesser space efficiency, with an associated performance trade-off. Figure 1 shows the strong performance preservation of using reduced-dimension linear heads. Halving the token dimension ($128\to 64$) only causes its nDCG@10 to drop by 0.01 (1.59%). We unfortunately find that MRL’s weight-tying efficient variant (MRL-E), where losses are computed on truncations of the same token vector does not preserve performance well.
@@ 5 Pair Training
To leverage an abundance of text pairs with varying richness of semantic structure, we draw inspiration from common practices in single-vector embedding model training and begin with training on these text pairs, focusing on optimizing the embedding model’s performance on general semantic similarity tasks. This weakly-supervised stage is in contrast to previous ColBERT works, which typically starts with the triplet training on 32-way or 64-way retrieval triplets, consisting of a query, a positive passage, and multiple mined negatives.
@@ 6.1 Data Composition
Our triplet dataset consists of 1) high-quality, human-annotated research datasets such as MSMARCO, DuReader, and MIRACL (Bajaj et al., 2016; He et al., 2018; Zhang et al., 2023b) with seven mined hard negatives per query and 2) high-quality datasets like MSMARCO and NQ translated from English into Chinese, French, German, Japanese, Russian and Spanish, following our previous work Mohr et al. (2024) and 3) synthetically generated datasets to address common failure modes of dense vector models such as negation and to cover niche domains like legal IR.
@@ 8.1 Efficient Evaluation
Due to the compute and time costs of indexing corpora containing tens of millions of documents, evaluating every model checkpoint and ablation on every task is not feasible. Therefore, we follow recent works (Clavié, 2024; Merrick et al., 2024) by comparing models’ quality on smaller sampled-corpus versions of HotpotQA, NQ, MS MARCO, and MIRACL (Chinese, French, German, Japanese, Spanish). These sampled corpora are constructed by combining the top 250 BM25-retrieved [^2] We use the standard pre-built Lucene indices in Pyserini (Lin et al., 2021) for MIRACL found at https://github.com/castorini/pyserini, and use BM25s (Lù, 2024) for BEIR. passages with all judged passages. We observe good agreement between the sampled-corpus evaluation scores and the full-fidelity ones when used to make binary or ranking-based model comparisons, but we leave a more rigorous analysis of this observation to future work.

### v4

@@ Abstract
Multi-vector dense models, such as ColBERT, have proven highly effective in information retrieval. ColBERT’s late interaction scoring approximates the joint query-document attention seen in cross-encoders while maintaining inference efficiency closer to traditional dense retrieval models, thanks to its bi-encoder architecture and recent optimizations in indexing and search. In this work we propose a number of incremental improvements to the ColBERT model architecture and training pipeline, using methods shown to work in the more mature single-vector embedding model training paradigm, particularly those that apply to heterogeneous multilingual data or boost efficiency with little tradeoff. Our new model, Jina-ColBERT-v2, demonstrates strong performance across a range of English and multilingual retrieval tasks.
@@ 1 Introduction
Neural retrieval has gained popularity in recent years following the arrival of capable pre-trained language models (PLMs) (Devlin et al., 2019; Liu et al., 2019; Clark et al., 2020). Two types of approaches have been employed to apply PLMs to retrieval. Sparse neural retrieval systems, such as SPLADE (Formal et al., 2021), represent texts as weighted bags of words that are interpreted as sparse high-dimensional vectors for maximum inner product search (MIPS). Dense retrievers similarly encode queries and documents as dense vectors, capturing relevance signals through spatial relationships extending beyond exact term matching.
Most dense retrievers encode a query or document as a single vector, commonly the result of mean-pooling or the [CLS]-embedding over the transformer’s final layer token embeddings. In contrast, recent multi-vector retrievers like ColBERT (Khattab and Zaharia, 2020) generalize this embedding process to maintain an embedding for each token, computing relevance scores as a function of the similarities of query and document tokens instead. To make the ColBERT usable in practice, the output dimensionality is restricted to be much smaller than the single-vector models. This approach has the benefit of remaining compatible with much of the vector similarity infrastructure that makes single-vector methods efficient, but requires more space to store even a smaller embedding per token and compute at inference time to aggregate token interactions into a single score. This late interaction over token embeddings achieves greater in-domain performance and tends to be more robust out-of-domain than single-vector similarity. While ColBERTv2 is trained only on English MSMARCO triplets (Bajaj et al., 2016) and has a monolingual BERT backbone, making it incapable of multilingual retrieval, some previous works extend the model to multilingual retrieval.
ColBERT-XM (Louis et al., 2024) does this by using parameter extensions for each additional language, and (Lawrie et al., 2023) trains solely on machine-translated English MSMARCO data to get effective heterogeneous multilingual performance. These approaches, however, come with trade-offs in terms of model usability and training data diversity. Other multilingual multi-vector models like BGE-M3 (Chen et al., 2024) produce extremely large token representations that limit their practical utility for first-stage retrieval.
@@
Training with diverse weakly-supervised data: We additionally pretrain our modified PLM with rotary position embedding and train on large-scale unlabeled text pairs from various corpus with a weakly-supervised single-vector contrastive objective. A second-stage of ColBERT finetuning with labeled triplet data and supervised distillation is used to further boost its performance.
General multilingual performance: We train with data from a variety of high- and low-resource languages using both labeled and unlabeled data, including human- and machine-translated training data, and show that this improves even out-of-domain multilingual performance.
Inference-agnostic efficiency: We introduce multiple sizes of linear projection heads, jointly trained using the non-weight tying variant of Matryoshka Representation Loss Kusupati et al. (2022), enabling the selection of token embedding size at inference time with minimal performance degradation. We demonstrate that reducing the embedding dimensionality in half from 128 to 64 yields only a minor performance tradeoff. Additionally, our flash-attention optimized backbone, Jina-XLM-RoBERTa  provides further free performance improvement during inference.
Our experimental results show competitive retrieval performance across both English and multilingual benchmarks. We also present controlled experiments demonstrating the benefits, or lack thereof, of the training modifications we consider in developing our training recipe.
@@ 2.3 Multilingual Retrieval
Owing to the quality of English-based pre-trained models (BERT) and annotated data (MSMARCO), many advances in neural retrieval have been applied first to the monolingual English setting Karpukhin et al. (2020); Xiong et al. (2020); Khattab and Zaharia (2020). Researchers, however, have also made advances in non-English capabilities.
On the modeling front, multilingual PLMs like mBERT (Devlin et al., 2019) and later XLM-RoBERTa (Conneau et al., 2020) have expanded pre-training to include text in up to 100 languages, including in cross-language contexts. For multilingual retrieval data, there are two approaches: natural and translated. Datasets like Mr-Tydi and MIRACL (Zhang et al., 2021, 2023b) are built from human-generated and annotated queries, whereas mMARCO (Bonifacio et al., 2022) is a collection of machine-translated copies of MSMARCO which inherit their judgments from the original dataset. The former method tends to be of higher quality and lacks the subtle distributional/idiomatic errors, dubbed "translationese", that the latter sometimes exhibits. Naturally, however, human generation costs more per example.
@@ 4.1 Backbone Improvements
We replace the absolute positional embeddings with rotary positional embeddings (RoPE, Su et al. (2023)), which are empirically understood to be better. They also have the advantage of supporting context lengths far longer than 512 tokens, although we do not explicitly focus on long-context in this work. To warm up its new positional embeddings, we continued pre-training the modified backbone with the same masked language modeling objective for 160,000 steps on the RefinedWeb dataset Penedo et al. (2023), a modern, high-quality corpus, under the masked language modeling objective. During this pre-training phase, we set the maximum sequence length to 8,192 tokens with a rotary base of 10,000 and employed whole-word-masking Devlin et al. (2019), masking out 30% of the tokens. We call this modified language model Jina-XLM-RoBERTa.
@@ 4.2 Multiple Linear Heads
To reduce index sizes, ColBERT includes a linear head that projects its token embeddings from the hidden dimension of its language model down to a lower dimension ($768\to 128$). As a notable exception, BGE-M3’s multi-vector retrieval does not take this step, keeping its token embeddings at a full 1024 dimensions.
We jointly train six linear heads with dimensions $d\in\{64,96,128,256,512,768\}$ using Matryoshka Representation Loss (MRL, Kusupati et al. (2022)). This allows users to choose greater or lesser space efficiency, with an associated performance trade-off. Figure 1 quantifies this tradeoff, showing the strong performance preservation of our reduced-dimension linear heads. Halving the token dimension ($128\to 64$) only causes its nDCG@10 to drop by 0.01 (1.59%). We unfortunately find that MRL’s weight-tying efficient variant (MRL-E), where losses are computed on truncations of the same token vector does not preserve performance well, which we hypothesize is a consequence of the already-low projected dimension of the original ColBERT formulation.
@@ 5 Pair Training
To leverage an abundance of text pairs with varying richness of semantic structure, we draw inspiration from common practices in single-vector embedding model training and begin with training on these text pairs, focusing on optimizing the embedding model’s performance on general semantic similarity and relatedness tasks. This weakly-supervised stage is in contrast to previous ColBERT works, which typically start directly from a PLM like BERT with triplet training on 32-way or 64-way retrieval triplets consisting of a query, a positive passage, and multiple mined negatives.
@@ 6.1 Data Composition
Our triplet dataset consists of 1) high-quality, human-annotated research datasets such as MSMARCO, DuReader, and MIRACL (Bajaj et al., 2016; He et al., 2018; Zhang et al., 2023b) with diversely mined negatives 2) high-quality datasets like MSMARCO and NQ translated from English into Chinese, French, German, Japanese, Russian and Spanish, following our previous work Mohr et al. (2024) and 3) synthetically generated datasets to address common failure modes of dense vector models such as negation and to cover niche domains like legal IR.
@@
Notably, owing to the limitations of our various sources of data, we train on triplets with only 7 negatives per example, in contrast to the 32- or 64-way triplets of ColBERTv2.
@@ 8.1 Efficient Evaluation
Due to the compute and time costs of indexing corpora containing tens of millions of documents, evaluating every model checkpoint and ablation on every task is not feasible. Therefore, we follow recent works (Clavié, 2024; Merrick et al., 2024) by comparing models’ quality on smaller sampled-corpus versions of HotpotQA, NQ, MS MARCO, and MIRACL (Chinese, French, German, Japanese, Spanish). These sampled corpora are constructed by combining the top 250 BM25-retrieved [^2] We use the standard pre-built Lucene indices in Pyserini (Lin et al., 2021) for MIRACL found at https://github.com/castorini/pyserini, and use BM25s (Lù, 2024) for BEIR. passages with all judged passages. We observe good agreement between the sampled-corpus evaluation scores and the full-fidelity ones when used to make binary or ranking-based model comparisons, but we leave a more rigorous analysis of this observation to future work. We only use the sampled corpora for ablation studies. For the final model, we evaluate on the full version of every dataset.